OpenStudy (samigupta8):

The value of summation of (n+1)^2/7^n is n starts from 0 to infinity...

OpenStudy (samigupta8):

My ans is 343/180 ...bt there is no option as such

OpenStudy (ben00):

I am not in high school.

OpenStudy (kainui):

Is this your sum? \[\sum_{n=0}^\infty \frac{(n+1)^2}{7^n}\]

OpenStudy (samigupta8):

Yep !

ganeshie8 (ganeshie8):

\[ \dfrac{1}{1-x}=\sum\limits_{n}x^n\] \[ \dfrac{1}{(1-x)^2}=\sum\limits_{n}nx^{n-1}\] \[ \dfrac{x}{(1-x)^2}=\sum\limits_{n}nx^{n}\] differentiate again

OpenStudy (kainui):

Since your sum looks a lot like the geometric series -- something we know exactly how to evaluate -- it's a good idea to see if you can recreate your series from the geometric series. So let's start here: \[\frac{1}{1-x} = \sum_{n=0}^\infty x^n\] Looks like we can plug in \(x = \tfrac{1}{7}\) and that is starting to look close, but we we want some \((n+1)\) factors, like what ganeshie is doing. :P

OpenStudy (samigupta8):

Well i did the same !! @kainui there will be two agp's solving them i got this result

OpenStudy (samigupta8):

First is dividing by 7 and subtract it from S Then again divide the term 6S/7 by 7 and subtract it from 6S/7 Finally we will have a gp of ratio 2/7

OpenStudy (kainui):

Well I guess I'll work it out and show my work so we can see what happened. \[\frac{1}{1-x} = \sum_{n=0}^\infty x^n\] Multiply both sides by x and then take the derivative: \[ \frac{x}{1-x} = \sum_{n=0}^\infty x^{n+1}\] \[\frac{d}{dx} \left( \frac{x}{1-x} \right) = \sum_{n=0}^\infty (n+1)x^{n}\] \[ \frac{1}{(1-x)^2} = \sum_{n=0}^\infty (n+1)x^{n}\] Repeat: \[ \frac{d}{dx} \left( \frac{x}{(1-x)^2} \right) = \sum_{n=0}^\infty (n+1)^2x^{n}\] \[ \frac{1+x}{(1-x)^3} = \sum_{n=0}^\infty (n+1)^2x^{n}\] Now plug in \(x = \tfrac{1}{7}\) and hope it works out: \[\frac{49}{27} = \sum_{n=0}^\infty (n+1)^27^{-n}\]

ganeshie8 (ganeshie8):

You want to work it w/o derivatives is it ?

OpenStudy (ben00):

Can that be in a new question?

OpenStudy (samigupta8):

Won't we be able to do if we try to solve the two agp's

OpenStudy (kainui):

I dunno, we can probably solve this some other way as well this just happened to be the way I saw to solve it first, I don't know what agp stands for sorry :X

OpenStudy (ben00):

Can I have some medals? :3

OpenStudy (samigupta8):

Arithmetic geometric progression

ganeshie8 (ganeshie8):

\[\begin{align}S &= \sum\limits_{n} \dfrac{n^2}{7^n} \\~\\\dfrac{1}{7}S &=\sum\limits_{n} \dfrac{n^2}{7^{n+1}}\\~\\ \dfrac{6}{7}S &=\sum\limits_{n}\dfrac{2n+1}{7^n} \end{align}\]

OpenStudy (samigupta8):

Correct again we will divide this by 7 and subtract it from 6s/7

OpenStudy (samigupta8):

Okk..i got the ans. .!!!

OpenStudy (samigupta8):

Sorry calculation mistake was being done by me...sorry guys....